OpenStudy (anonymous):

Would x^1/3 times x^1/3 be x^2/6?

OpenStudy (anonymous):

possible answers, x^1/3 x^2/6 x^2/3 x*1/9

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@confluxepic

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (radar):

When multiplying common bases, you simply "add" the exponents.

OpenStudy (radar):

Soooooo what is....... 1/3 + 1/3 = ??????

OpenStudy (anonymous):

2/6

OpenStudy (zephyr141):

no. 1/2 +1/2 is not 2/4 it's 1. neither is 1/5 +1/5 = 2/10 . those are wrong. 1/2 +1/2 =1 and 1/5+1/5=2/5 and 1/3+1/3=2/3